The manager of a fast food restaurant with a drive up window is interested in how long people have to wait between the time they order and the time their order is filled. The assistant manager randomly samples 285 customers and records their wait time. The mean wait time was 340 seconds with a standard deviation of 73. What is the upper limit of the 99% confidence interval for mean wait time? Round up to the nearest whole seconds.
Answer #1


Solution :

Given that,

= 340

s = 73

n = 285

Degrees of freedom = df = n - 1 = 285- 1 = 284

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

t /2,df = t0.005,284 = 2.593

Margin of error = E = t/2,df * (s /n)

= 2.593 * (73 / 285)

= 11.21

Margin of error = 11.21

The 99% confidence interval estimate of the population mean is,

< + E

340 +11.21

The upper limit = 351.21